OpenStudy (anonymous):

HELP! given: abc, ab is congruent to ac. d is a point between b and c prove: ab>ad

OpenStudy (mathmale):

Hey, Em: It would help me, and perhaps help you as well, if you'd define abc. I suspect you mean the triangle abc, but it's possible that one could correctly identify abc as a straight line going through 3 points. See the ambiguity?

OpenStudy (anonymous):

its a triangle

OpenStudy (anonymous):

@mathmale

OpenStudy (ranga):



OpenStudy (ranga):

From (2) we can see ADB will be highest when x is lowest and ADB will be lowest when x is highest. From (1) we know the lowest and highest value of x. Let us find the lowest value of ADB by putting in the highest value of x from (1) into (2) Lowest ADB = 180 - B - (180 - 2B) = 180 - B - 180 + 2B = B. Lowest ADB is B and that happens when D coincides with C. But we are told that D is a point in between B and C. So ADB will ALWAYS be greater than B. Therefore, AB will always be greater than AD in the triangle ABD.

OpenStudy (anonymous):

i almost understand this but on my chart c is were b is and vise versa

OpenStudy (ranga):

Does not matter. It is an isosceles triangle and so angle B = angle C and it does not matter which is B and which is C.

OpenStudy (ranga):

But if there was a diagram that went with the problem and you want the answer to be consistent with that chart you should have drawn the chart so we could have answered accordingly. But you can get the idea of what we are trying to do here and modify accordingly.

OpenStudy (anonymous):

okay i just didnt know how to insert the chart with the problem

OpenStudy (ranga):

You can click on the Draw button below the Reply box and draw a diagram like I did it. Or you can take a "print screen" of the image on your screen and upload the picture. Or you can take a picture of the page and post it here, etc.

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

 The same answer applies to this new drawing. You are welcome.